OpenStudy (shelby1290):

Which expression is equivalent to -7(x-2) + 5(3-x)-4x? a) -16x +1 b) -16x+29 c) -2x + 1 d) -12x + 13

OpenStudy (shelby1290):

I got D as my answer, am i correct?

OpenStudy (shelby1290):

-7x -2 + 15 - x - 4x -7x-x-4x-2 +15 -12x=-2+15 -12x=13

OpenStudy (anonymous):

-7x + 14 +15 - 5x -4x -16x + 29

OpenStudy (shelby1290):

@jualbafer how'd you get 14?

OpenStudy (shelby1290):

i'm confused